# Instance-adaptive Zero-shot Chain-of-Thought Prompting

### v2

@@ Abstract
Zero-shot Chain-of-Thought (CoT) prompting emerges as a simple and effective strategy for enhancing the performance of large language models (LLMs) in real-world reasoning tasks. Nonetheless, the efficacy of a singular, task-level prompt uniformly applied across the whole of instances is inherently limited since one prompt cannot be a good partner for all, a more appropriate approach should consider the interaction between the prompt and each instance meticulously. This work introduces an instance-adaptive prompting algorithm as an alternative zero-shot CoT reasoning scheme by adaptively differentiating good and bad prompts. Concretely, we first employ analysis on LLMs through the lens of information flow to detect the mechanism under zero-shot CoT reasoning, in which we discover that information flows from question to prompt and question to rationale jointly influence the reasoning results most. We notice that a better zero-shot CoT reasoning needs the prompt to obtain semantic information from the question then the rationale aggregates sufficient information from the question directly and via the prompt indirectly. On the contrary, lacking any of those would probably lead to a bad one. Stem from that, we further propose an instance-adaptive prompting strategy (IAP) for zero-shot CoT reasoning. Experiments conducted with LLaMA-2, LLaMA-3, and Qwen on math, logic, and commonsense reasoning tasks (e.g., GSM8K, MMLU, Causal Judgement) obtain consistent improvement, demonstrating that the instance-adaptive zero-shot CoT prompting performs better than other task-level methods with some curated prompts or sophisticated procedures, showing the significance of our findings in the zero-shot CoT reasoning mechanism.
@@ 1 Introduction
Inspired by the above findings, we further propose an instance-adaptive prompting strategy (IAP) for zero-shot CoT reasoning. Given a list of prompts in distinct styles, we try to recognize good ones that elicit LLMs to reason toward the correct answer while avoiding bad CoT reasoning, referring to the analytical results. We conduct comprehensive experiments with IAP and existing methods with multiple LLMs on various tasks. Experimental results show that the IAP can consistently improve the overall performance of LLMs such as LLaMA-2 13B, LLaMA-3 8B, and Qwen 14B on kinds of reasoning tasks including math, logic, and commonsense reasoning. Specifically, the IAP strategy achieves a 2%-4% accuracy enhancement across tasks and models compared to the optimal task-level prompt. Our contributions can be summarized as follows:
@@ 2.1 Preliminary analysis
We define the reasoning that produces the right answer to a given question as good reasoning; otherwise as bad reasoning. Given various information interactions happen among reasoning steps during the model inference, choosing which step (i.e., output token) to explore is critical. Despite the most popular practice being the last step for the In-Context Learning [19, 21], there are distinct circumstances in the CoT reasoning [32], our investigation shows that not all the final answers appear at the reasoning last step. To eliminate the effect of distinct LLM generation styles as much as possible, we adopt uniformly the answer generation step for all tasks as our observation time, concretely, we implement that with several regular expressions to recognize the answer step during model inference. More details are in Appendix A.2.
@@
Since "Let’s think step by step." and "Don’t think. Just feel." are two representative good and bad zero-shot prompts on task-level in  [2], we select them as our test prompts. To inspect saliency scores of good reasonings and bad ones, we randomly pick two pairs of good-bad reasoning instances under two prompts and visualize the saliency scores inside them in Figure 2, each subfigure depicts the mean of saliency matrices of all layers and all heads, i.e., $I=\frac{1}{LH}\sum_{\ell=1}^{L}\sum_{h=1}^{H}I^{(\ell,h)}$ where $L$ and $H$ are the numbers of layer and head. As mentioned earlier, we emphasize the question, prompt, and rationale during reasoning. In Figure 2(a), tokens from the first to the last of the prompt collect information from the question tokens evidently, and some tokens especially those near the answer in the rationale aggregate information from the question and prompt tokens evidently, either. In Figure 2(b), things start to change, it seems that the prompt tokens fail to gather information from question tokens, not sufficiently at least, and tokens in the rationale are unable to gain much information from the question or the prompt.
@@ 3 Adaptive Instance-level Zero-shot CoT
Based on the above findings, we believe that a prompt with saliency scores surpassing the corresponding threshold is considered a good prompt for a given question, consequently mitigating the need to explore further prompts. This process terminates upon either identifying an optimal prompt or traversing all candidates.
@@ 4.1 Implementation
Models. We test IAP and comparison methods on LLaMA-3 8B [42], LLaMA-2 13B [41], and Qwen 14B [43] since they are popular Transformer decoder-only LLMs, which is convenient for exploiting and analyzing inside architectures. We set the generation mode to greedy-decoding to minimize irrelevant confounders during the model inference to ensure the answers to fixed questions under the same model and prompt, and all the experiments are run on an 8x NVIDIA A100 GPU server.
Baselines. OPPR [22] takes the LLM as an optimizer to update a zeros-hot CoT prompt iteratively and produce corresponding optimized prompts for copious tasks. Self-Discover [24] selects relevant atomic reasoning modules (e.g., decomposing problems, critical thinking) for a given task, then adapts and customizes those modules to fit the task. These two frameworks aim to search for an appropriate prompt, which is essentially similar to our purpose, so we choose them as comparisons with the IAP to observe the performance difference between instance-level and task-level zero-shot CoT prompting.
@@
Table 1: Zero-shot CoT results with LLaMA-3 8B and Qwen 14B under various prompts, the results of LLaMA-2 13B is in Appendix A.1. Each column stands for a group of task categories, T-Obj. is for Tracking Shuffled Objects which are from the BBH. The "Optimizer-generated prompt" refers to the prompts for each task generated with the algorithm in [22].
@@ 4.2 Results
Prompts steer these LLMs to achieve different results in multiple tasks, and no single prompt can get an overwhelming performance on all datasets, which makes our research on the mechanism of zero-shot CoT valuable. Table 1 shows the zero-shot CoT reasoning results with LLaMA-3 8B and Qwen 14B and various prompts on 3 reasoning tasks, we put the results of LLaMA-2 13B in the Appendix since LLaMA models share a quite similar architecture.
Math reasoning. Compared with these prompts, IAP-mv improves the LLaMA-3 and Qwen’s accuracy on GSM8K from 64.52%, 60.50% to 66.34%, 62.81% respectively. On SVAMP, IAP-mv obtains a 1.33% improvement on both two models compared to the task-level-optimal prompt. It is worth noting that OPPR and Self-discover are unstable with different LLMs and datasets, indicating the unstable characteristics of task-level prompting. Results on these two math reasoning datasets demonstrate the IAP can benefit the math reasoning task.
Logic reasoning. For Causal Judgement, IAP-ss and IAP-mv enhance the accuracy of the task-level optimal prompt and IAP-mv outperforms OPPR, which is optimized by numerous iterations. For Tracking shuffle Objects, IAP-mv performs well with Qwen while achieving a sub-optimal accuracy with LLaMA-3, IAP-mv still obtains strong results, improving 2.4% and 2.3% with LLaMA-3 and Qwen, respectively.
Commonsense reasoning. On CSQA, the IAP improves the accuracy of the former best for 3.44% with LLaMA-3, and 2.45% with Qwen. On MMLU, LLaMA-3 and Qwen obtain improvement to a large margin, either. We note that improving IAP-mv and IAP-ss on commonsense reasoning is quite salient, demonstrating the effectiveness of the saliency score-based prompting strategies.
Table 2: Accuracy (%) of Consistency and Complementary prompts with IAP-mv on 3 tasks with LLaMA-3 8B and Qwen 14B. The results of LLaMA-2 13B are at Appendix A.1.
@@
Figure 6: Efficiency comparison with LLaMA-3 8B on the Tracking Shuffle Objects, 0-shot denotes the best task-level prompt.
@@ 4.3 Ablation studies
The success of zero-shot prompting for CoT reasoning lies in the semantic information within those prompts, when the LLM receives a prompt, it would generate rationales by obeying the meaning of the prompt as much as possible. According to semantics,  [2] categorizes these zero-shot CoT prompts into 3 types: instructive, misleading, and irrelevant, and we further define that prompts in the same category are consistent, or otherwise they are complementary. Our 9 prompts can be divided into 3 consistency groups, but the irrelevant group contains only one prompt, thus we evaluate the complementary on the other two. For the complementary groups, we build them two-by-two. Table 2 depicts the performance of each group. We employ IAP-mv since it manifests a stronger capability in taking advantage of multiple prompts.
@@
As we mentioned in Section 3, the IAP-mv trades efficiency for performance, and IAP-ss emphasizes efficiency. We introduce the reasoning time (seconds) for each iteration complete as the metric to measure the efficiency and conduct time-consuming experiments under the same setting to show the cost of IAP-mv, IAP-ss, and Self-Discover on the Tracking Shuffle Objects dataset, results are shown in Figure 6. All these strategies increase the computation cost to a certain degree, while IAP-ss may bring accuracy decreases than the task-level optimal prompt, it beats Self-discover. Though IAP-mv is the most time-consuming, it can improve performance, therefore, the two IAP strategies can be employed as trade-offs in different demand prioritization applications.
@@ 5 Related Work
Advanced prompting approaches such as SelfzCoT [30] and Meta-prompting [31] showcase the evolutionary trajectory of prompting, which generates semantic and code prompts through a root prompt to obtain precise answers, while Meta-prompting deconstructs complex tasks into simpler sub-tasks, each addressed by specialized models to foster inter-model communication and apply intricate reasoning. Lastly, methodologies like OPRO [22] and the innovative concept of evolutionary prompting [23] aim to recursively optimize CoT prompts and generate varied prompts through mutations and crossovers, respectively. Self-discover [24] selects relevant atomic reasoning modules (e.g., breaking down problems, critical thinking) for a given task, then adapts and customizes those modules to fit the task. Implement the customized reasoning structure when solving task instances. These workarounds significantly contribute to the development of zero-shot CoT prompts that guide LLMs toward more accurate problem framing, intermediate reasoning, and final answers.
@@ 6 Conclusion
In this paper, we aim to delve into the mechanism of LLMs in zero-shot CoT reasoning from the perspective of information flow to understand what happened during this process, and we find stronger saliency scores within question-to-prompt and question-to-rationale can lead to better LLM reasoning. To investigate these phenomena in a nuanced manner, we go deep into the Transformer layers and attention heads in the LLM and find the front of the middle and final heads in shallow layers carry more information during information flows. Inspired by that, we present an instance-adaptive zero-shot prompting strategy for better CoT reasoning. To demonstrate our findings, we conduct comprehensive experiments on several LLMs and tasks, and the results show our proposed strategies can improve the performance of LLMs on all candidate prompts, highlighting our interpretation of zero-shot CoT in the view of information flow.
@@
In this work, we select the answer step as the key step to investigate the saliency scores, even in most instances it can be located well, and some irregular answers can not be identified precisely, such a factor may affect the accuracy of our analysis. Despite our research in this paper providing insight into understanding the underlying workflow of zero-shot CoT reasoning, it cannot be the only interpretation, and we believe there must be better means to explain that.
@@ Appendix A Appendix
A.1 IAP & Baseline Experiments on LLaMA-2 13B
Table 3 and Table 4 are supplementary for Table 1 and Table 1, the results here basically are coincidence with the Experiment section in the main body. However, the IAP can not obtain consistent enhancement, which may caused by the irregular output formats of LLaMA-2.
Table 3: Comparison with baseline methods using accuracies on MultiArith and GSM8K. LLaMA-2 13B is used as the model if not specified. We report the results when no shot-coT is applied. Each column stands for a group of task categories, T-Obj. is for Tracking Objects which are from the BBH. The "Optimizer-generated prompt" refers to the prompts for each task generated with the algorithm in [22].

### v3

@@ Abstract
Zero-shot Chain-of-Thought (CoT) prompting emerges as a simple and effective strategy for enhancing the performance of large language models (LLMs) in real-world reasoning tasks. Nonetheless, the efficacy of a singular, task-level prompt uniformly applied across the whole of instances is inherently limited since one prompt cannot be a good partner for all, a more appropriate approach should consider the interaction between the prompt and each instance meticulously. This work introduces an instance-adaptive prompting algorithm as an alternative zero-shot CoT reasoning scheme by adaptively differentiating good and bad prompts. Concretely, we first employ analysis on LLMs through the lens of information flow to detect the mechanism under zero-shot CoT reasoning, in which we discover that information flows from question to prompt and question to rationale jointly influence the reasoning results most. We notice that a better zero-shot CoT reasoning needs the prompt to obtain semantic information from the question, and then the rationale aggregates sufficient information from the question directly and via the prompt indirectly. On the contrary, lacking any of those would probably lead to a bad one. Stem from that, we further propose an instance-adaptive prompting strategy (IAP) for zero-shot CoT reasoning. Experiments conducted with LLaMA-2, LLaMA-3, and Qwen on math, logic, and commonsense reasoning tasks (e.g., GSM8K, MMLU, Causal Judgement) obtain consistent improvement, demonstrating that the instance-adaptive zero-shot CoT prompting performs better than other task-level methods with some curated prompts or sophisticated procedures, showing the significance of our findings in the zero-shot CoT reasoning mechanism.
@@ 1 Introduction
Inspired by the above findings, we further propose an instance-adaptive prompting strategy (IAP) for zero-shot CoT reasoning. Given a list of prompts in distinct styles, we try to recognize good ones that elicit LLMs to reason toward the correct answer while avoiding bad CoT reasoning, referring to the analytical results. We conduct comprehensive experiments with IAP and existing methods with multiple LLMs on various tasks. Experimental results show that the IAP can consistently improve the overall performance of LLMs such as LLaMA-2-13B-Chat, LLaMA-3-8B-Instruct, and Qwen-14B-Chat on kinds of reasoning tasks including math, logic, and commonsense reasoning. Specifically, the IAP strategy achieves a 2%-4% accuracy enhancement across tasks and models compared to the optimal task-level prompt. Our contributions can be summarized as follows:
@@ 2.1 Preliminary analysis
We define the reasoning that produces the right answer to a given question as good reasoning, otherwise as bad reasoning. Given various information interactions happen among reasoning steps during the model inference, choosing which step (i.e., output token) to explore is critical. Despite the most popular practice being the last step for the In-Context Learning [19, 21], there are distinct circumstances in the CoT reasoning [32], our investigation shows that not all the final answers appear at the reasoning last step. To eliminate the effect of distinct LLM generation styles as much as possible, we adopt uniformly the answer generation step for all tasks as our observation time, concretely, we implement that with several regular expressions to recognize the answer step during model inference. More details are in Appendix A.2.
@@
Since "Let’s think step by step." and "Don’t think. Just feel." are two representative good and bad zero-shot prompts on task-level in  [2], we select them as our test prompts. We explore the saliency score with Qwen-14B [43] on GSM8K [36] and maintain consistency in the following analysis, and we also put similar visual analysis on other models and datasets in the Appendix. To inspect saliency scores of good reasonings and bad ones, we randomly pick two pairs of good-bad reasoning instances under two prompts and visualize the saliency scores inside them in Figure 2, each subfigure depicts the mean of saliency matrices of all layers and all heads, i.e., $I=\frac{1}{LH}\sum_{\ell=1}^{L}\sum_{h=1}^{H}I^{(\ell,h)}$ where $L$ and $H$ are the numbers of layer and head. As mentioned earlier, we emphasize the question, prompt, and rationale during reasoning. In Figure 2(a), tokens from the first to the last of the prompt collect information from the question tokens evidently, and some tokens especially those near the answer in the rationale aggregate information from the question and prompt tokens evidently, either. In Figure 2(b), things start to change, it seems that the prompt tokens fail to gather information from question tokens, not sufficiently at least, and tokens in the rationale are unable to gain much information from the question or the prompt.
@@ 3 Adaptive Instance-level Zero-shot CoT
Based on the above findings, we believe that a prompt with saliency scores surpassing the corresponding threshold is considered a good prompt for a given question, consequently mitigating the need to explore further prompts. Given the training data, we can search for a This process terminates upon either identifying an optimal prompt or traversing all candidates.
@@ 4.1 Implementation
Models. We test IAP and comparison methods on LLaMA-3-8B-Instruct [42], LLaMA-3-70B-Instruct , LLaMA-2-13B-Chat [41], and Qwen-14B-Chat [43] since they are popular Transformer decoder-only LLMs, which is convenient for exploiting and analyzing inside architectures. For the IAP-ss, we obtain threshold values w.r.t distinct LLMs on different datasets, we compute the overall synthesized scores as defined in Eq 4. We set the generation mode to greedy-decoding to minimize irrelevant confounders during the model inference to ensure the answers to fixed questions under the same model and prompt, and all the experiments are run on an 8x NVIDIA A100 GPU server.
Baselines. Answer majority vote (AMV) is a simple method implemented by choosing the most popular result of all prompt candidates for a given question as its final answer. OPPR [22] takes the LLM as an optimizer to update a zeros-hot CoT prompt iteratively and produce corresponding optimized prompts for copious tasks. Self-Discover [24] selects relevant atomic reasoning modules (e.g., decomposing problems, critical thinking) for a given task, then adapts and customizes those modules to fit the task. These two frameworks aim to search for an appropriate prompt, similar to our purpose. We choose them as comparisons with the IAP to observe the performance difference between instance-level and task-level zero-shot CoT prompting.
@@
Table 1: Zero-shot CoT results with LLaMA-3-8B-Instruct and Qwen-14B-Chat under various prompts, the results of LLaMA-3-70B-Instruct and LLaMA-2-13B-Chat is in Appendix A.1. Each column stands for a group of task categories, T-Obj. is for Tracking Shuffled Objects which are from the BBH. The "Optimizer-generated prompt" refers to the prompts for each task generated with the algorithm in [22].
@@ 4.2 Results
Prompts steer these LLMs to achieve different results in multiple tasks, and no single prompt can get an overwhelming performance on all datasets, which makes our research on the mechanism of zero-shot CoT valuable. Table 1 shows the zero-shot CoT reasoning results with LLaMA-3-8B-Instruct and Qwen-14B-Chat and various prompts on 3 reasoning tasks, we put the results of LLaMA-2-13B-Chat and larger LLaMA-3 70B in the Appendix since LLaMA models share a quite similar architecture.
Math reasoning. Compared with these prompts, IAP-mv improves the LLaMA-3 and Qwen’s accuracy on GSM8K from 64.52%, 60.50% to 66.34%, 62.81% respectively. On SVAMP, IAP-mv obtains a 1.33% improvement on both models compared to the task-level-optimal prompt. It is worth noting that OPPR and Self-discover are unstable with different LLMs and datasets, indicating the unstable characteristics of task-level prompting. Results on these two math reasoning datasets demonstrate the IAP can benefit the math reasoning task.
Logic reasoning. For Causal Judgement, IAP-ss and IAP-mv enhance the accuracy of the task-level optimal prompt and IAP-mv outperforms OPPR, which is optimized by numerous iterations. For Tracking shuffle Objects, IAP-mv performs well with Qwen while achieving a sub-optimal accuracy with LLaMA-3, IAP-mv still obtains strong results, improving 2.4% and 2.3% with LLaMA-3 and Qwen, separately.
Commonsense reasoning. On CSQA, the IAP improves the accuracy of the former best for 3.44% with LLaMA-3, and 2.45% with Qwen. On MMLU, LLaMA-3 and Qwen obtain improvement to a large margin, either. We note that improving IAP-mv and IAP-ss on commonsense reasoning is more salient than the other two reasoning tasks, demonstrating the effectiveness of the saliency score-based prompting strategies.
Table 2: Accuracy (%) of Consistency and Complementary prompts with IAP-mv on 3 tasks with LLaMA-3-8B-Instruct and Qwen-14B-Chat. The results of LLaMA-2-13B-Chat are at Appendix A.1.
@@
Apart from the above comparison, we can also observe that the answer majority vote among all prompts performs poorly in some tasks, in contrast, AMV (#1-7) can enhance it by eliminating misleading prompts. Such results indicate the instability of the AMV, whose performance can be affected by prompt candidates. Our IAP-mv outperforms it by a large margin, demonstrating that most prompts can lead the LLM to generate wrong answers for a given question, reaching only a few correct answers, i.e., such methods cannot recognize good or bad reasoning. Our IAP-mv can handle that with the analysis for information flow in reasoning, i.e., IAP-mv can differentiate good and bad reasoning, validating the effectiveness of our proposed strategy.
Figure 6: Efficiency comparison with LLaMA-3-8B-Instruct on the Tracking Shuffle Objects, 0-shot denotes the best task-level prompt.
@@ 4.3 Ablation Studies
The success of zero-shot prompting for CoT reasoning lies in the semantic information within those prompts, when the LLM receives a prompt, it would generate rationales by obeying the meaning of the prompt as much as possible. According to semantics,  [2] categorizes these zero-shot CoT prompts into 3 types: instructive, misleading, and irrelevant, and we further define that prompts in the same category are consistent, or otherwise they are complementary. To detect which type of prompt combination contributed to the performance, we divide the 9 prompts into 3 consistency groups, but the irrelevant group contains only one prompt, thus we evaluate the complementary on the other two. For the complementary groups, we build them two-by-two. Table 2 depicts the performance of each group. We employ IAP-mv since it manifests a stronger capability in harnessing multiple prompts. we can observe that each pair of combinations can improve the performance, and instructive and irrelevant combinations achieve better outcomes than others, which comes from the base performance of instructive prompts.
@@
The order and number of prompt candidates are critical for the accuracy and efficacy of IAP-ss, in this paper, we adopt the #1-9 order to conduct IAP, and we also tried other settings. In Table 3., #9 is the worst task-level prompt, and #6 is the best task-level prompt, achieving the highest accuracy among all the prompt candidates while consuming the least time.
Table 3: Accuracy and inference time (s) with different prompt orders and numbers of LLaMA-3-8B-Instruct on SVAMP.
  Order Acc Time
#9 39.67 2860
#6 76.00 2657
#9, 8, 5, 4, 3 63.66 3870
#6, 1, 2, 7, 3 76.66 5216
#1, 2, 3, 4, 5, 6, 7, 8, 9 77.33 4751
The prompt order of the 3rd row is accuracy-decreased on SVAMP, the 4th row is accuracy-increased, and the last row is our default setting, which obtains the best performance. This table shows that IAP-ss can cost less time with fewer prompt candidates but may obtain limited results, however, even fewer improper candidates could take a lot of computing time. Therefore, the time cost of IAP-ss is not a major issue if prompt candidates are in an appropriate order. As we mentioned in Section 3, the IAP-mv trades efficiency for performance, and IAP-ss emphasizes efficiency. We introduce the reasoning time (seconds) for each iteration complete as the metric to measure the efficiency and conduct time-consuming experiments under the same setting to show the cost of IAP-mv, IAP-ss, and Self-Discover on the Tracking Shuffle Objects dataset, results are shown in Figure 6. All these strategies increase the computation cost to a certain degree, while IAP-ss may bring accuracy decreases than the task-level optimal prompt, it beats Self-discover. Though IAP-mv is the most time-consuming, it can improve performance, therefore, the two IAP strategies can be employed as trade-offs in different demand prioritization applications.
@@ 5 Related Work
Advanced prompting approaches such as SelfzCoT [30] and Meta-prompting [31] showcase the evolutionary trajectory of prompting, which generates semantic and code prompts through a root prompt to obtain precise answers, while Meta-prompting deconstructs complex tasks into simpler sub-tasks, each addressed by specialized models to foster inter-model communication and apply intricate reasoning. Lastly, methodologies like OPRO [22] and the innovative concept of evolutionary prompting [23] aim to recursively optimize CoT prompts and generate varied prompts through mutations and crossovers. Self-discover [24] selects relevant atomic reasoning modules (e.g., breaking down problems, critical thinking) for a given task, then adapts and customizes those modules to fit the task. Implement the customized reasoning structure when solving task instances. These workarounds significantly contribute to developing zero-shot CoT prompts that guide LLMs toward more accurate problem framing, intermediate reasoning, and final answers.
@@ 6 Conclusion
In this paper, we aim to delve into the mechanism of LLMs in zero-shot CoT reasoning from the perspective of information flow to understand what happened during this process, and we find stronger saliency scores within question-to-prompt and question-to-rationale can lead to better LLM reasoning. To investigate these phenomena nuancedly, we go deep into the Transformer layers and attention heads in the LLM and find the front of the middle and final heads in shallow layers carry more information during information flows. Inspired by that, we present an instance-adaptive zero-shot prompting strategy for better CoT reasoning. To demonstrate our findings, we conduct comprehensive experiments on several LLMs and tasks, and the results show our proposed strategies can improve the performance of LLMs on all candidate prompts, highlighting our interpretation of zero-shot CoT in the view of information flow.
@@
In this work, we select the answer step as the key step to investigate and visualize the saliency scores, even in most instances it can be located well, and some irregular answers can not be identified precisely, such a factor may affect the generality and accuracy of our analysis. Different LLMs may have distinct patterns under the zero-shot CoT reasoning, for example, our analysis and conclusion can not meet all models. Despite our research providing insight into understanding the underlying workflow of zero-shot CoT reasoning, it cannot be the only interpretation, and we believe there must be better means to explain that.
Acknowledgement
This work is supported in part by Alibaba Research Intern Program, the National Natural Science Foundation of China (No. 62272191), the International Science and Technology Cooperation Program of Jilin Province (No. 20240402067GH), and the Science and Technology Development Program of Jilin Province (No. 20220201153GX).
@@ Appendix A Appendix
A.1 IAP & Baseline Experiments on LLaMA-2-13B-Chat and LLaMA-3 70B
Table 4 and Table 6 are supplementary for Table 1 and Table 2, the results here basically are coincidence with the Experiment section in the main body. Table 5 shows the results of single prompts and IAP-mv to demonstrate the generality for a large LLM. However, the IAP did not obtain salient enhancement, which may caused by the irregular output formats of LLaMA-2.
Table 4: Zero-shot CoT results with LLaMA-2-13B-Chat under various prompts and other baselines.